solve 2/3 + 1/4x = -1/3
Answer by london maths tutor(243)

2/3 + 1/(4x) = -1/3
The common denominator = 3*(4x) = 12x
[(2*4x)+(1*3)]/12x = -1/3
[3*(8x+3)]=12x*(-1)
24x+9 = - 12x
24x+12x = -9
36x = -9
x = -9/36
x = -1/4